# Some Constructions on Quantum Principal Bundles

### v4

@@ 3. The Group Isomorphism
$=\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})\otimes\mathbbm{1})\widetilde{\beta}((\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathfrak{F}^{-1})\mathrm{qtrs}(\vartheta^{(2)}))$
[…]
$=(-1)^{\partial\vartheta^{(2)}\partial[\vartheta^{(4)}]_{1}}\,\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\;\mathfrak{F}([\vartheta^{(1)}]_{2})\otimes\underbrace{\epsilon(\vartheta^{(2)})\mathbbm{1}}_{0-\mathrm{degree}})\,\widetilde{\beta}((\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathfrak{F}^{-1})\mathrm{qtrs}(\vartheta^{(3)})).$
$=(-1)^{(\partial\vartheta^{(3)}+\partial\vartheta^{(2)})\partial[\vartheta^{(4)}]_{1}}\,\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})\otimes\underbrace{\vartheta^{(2)}S(\vartheta^{(3)}}_{0-\mathrm{degree}}))\,\widetilde{\beta}((\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathfrak{F}^{-1})\mathrm{qtrs}(\vartheta^{(4)}))$
$=(-1)^{(\partial\vartheta^{(3)}+\partial\vartheta^{(2)})\partial[\vartheta^{(4)}]_{1}}\,\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})\otimes\underbrace{\vartheta^{(2)}S(\vartheta^{(3)}}_{0-\mathrm{degree}}))\,([\vartheta^{(4)}]_{1}\otimes\mathbbm{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}^{-1}([\vartheta^{(4)}]_{2}))$
$=(-1)^{(\partial\vartheta^{(3)}+\partial\vartheta^{(2)})\partial[\vartheta^{(4)}]_{1}}\,\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})[\vartheta^{(4)}]_{1}\otimes\vartheta^{(2)}S(\vartheta^{(3)}))\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}^{-1}([\vartheta^{(4)}]_{2}))$
@@
$=\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})[\vartheta^{(2)}]_{1}\otimes\mathbbm{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}^{-1}([\vartheta^{(2)}]_{2})),$
which proves our claim. Since $\mathfrak{F}^{-1}$ is a left $\Omega^{\bullet}(B)$–module morphism, by our previous claim and Proposition 3.3 point $1$
@@
for every $\vartheta$in $\Gamma^{\wedge}$. In a similar way, we can prove that $\mathfrak{f}_{\mathfrak{F}^{-1}}\ast\mathfrak{f}_{\mathfrak{F}}=\mathbbm{1}\epsilon$ and hence $\mathfrak{f}_{\mathfrak{F}}$ is a convolution invertible map.
@@
We are going to prove that $\mathfrak{F}_{\mathfrak{f}}$ is a graded left $\Omega^{\bullet}(B)$–module isomorphism which satisfies equation (45). First of all, by equation (43) it is obvious that $\mathfrak{F}_{\mathfrak{f}}(\mathbbm{1})=\mathbbm{1}$. Secondly, taking $\mu$in $\Omega^{\bullet}(B)$ and $w$in $\Omega^{\bullet}(P)$ we have $\mathfrak{F}_{\mathfrak{f}}(\mu w)=m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes\mathfrak{f})(\mu\otimes\mathbbm{1})\Delta_{\Omega^{\bullet}(P)}(w)=\mu\,\mathfrak{F}_{f}(w).$ Thirdly, by equation (43)
@@
[…]
@@
Finally, notice that for all $w$in $\Omega^{\bullet}(P)$
@@
are mutually inverse. Notice that for all $w$in $\Omega^{\bullet}(P)$
@@
[…]
@@
On the other hand, for every $\vartheta$in $\Gamma^{\wedge}$
@@
for all $w$in $\Omega^{\bullet}(P)$. Therefore $\mathfrak{f}_{1}\ast\mathfrak{f}_{2}\xmapsto{\widehat{\Lambda}^{-1}}\mathfrak{F}_{\mathfrak{f}_{1}}\circ\mathfrak{F}_{\mathfrak{f}_{2}}$ and the proposition follows. ∎
@@ 4. The form of the maps $\{T^{\mathrm{L}}_{k}\}$
Let $G$ be a compact matrix Lie group, and let $\mathcal{G}$ be its associated quantum group [2]. If $\delta^{V}$ is a (unitary) irreducible $\mathcal{G}$–corepresentation, then it induces a $G$–representation on $V$. Indeed, by Theorem 1.1 we know that
@@
is a unitary and irreducible $G$–representation. The following proposition is another of our purposes.
@@
Let $G$ be a compact matrix Lie group and let $\pi:P\longrightarrow B$ be a classical principal $G$–bundle, where $P$ is the total space, $B$ is the base space and $\pi$ is the bundle projection. Assume $B$ are compact. If $\mathcal{T}$ is a complete set of mutually non–equivalent irreducible $\mathcal{G}$–corepresentations with $\delta^{\mathbb{C}}_{\mathrm{triv}}$in $\mathcal{T}$, then for every $\delta^{V}$in $\mathcal{T}$, there exists $\{T^{\mathrm{L}}_{k}\}^{d_{V}}_{k=1}$subseteq $\textsc{Mor}(\delta^{V},\Delta_{P})$, for some $d_{V}$in $\mathbb{N}$, such that equation (30) holds. Here
@@
is the pull–back of the right $G$–action on $P$. In this setting, $C^{\infty}_{\mathbb{C}}(P)$ denotes the space of $\mathbb{C}$–valued smooth functions of $P$ and the tensor product is taken to be the completed injective tensor product.
@@
Let $\delta^{V}$in $\mathcal{T}$ and consider its associated representation $\alpha^{V}$ with $n=\mathrm{dim}_{\mathbb{C}}(V)$. Since $\alpha^{V}$ is unitary, the associated vector bundle $\pi_{\alpha^{V}}:E^{V}\longrightarrow B$ is a Hermitian bundle.
For each $b$in $B$, let $(U_{b},\Phi_{b})$ be a local trivialization of the Hermitian bundle $\pi_{\alpha^{V}}:E^{V}\longrightarrow B$ around $b$ associated to a principal $G$–bundle local trivialization $(U_{b},\Psi_{b})$ of $\pi:P\longrightarrow B$. Then there exists a set of local sections $\{\hat{s}^{b}_{1},\ldots,\hat{s}^{b}_{1}\}$subseteq $\Gamma\bigl{(}\pi_{{\alpha^{V}}}^{-1}(U_{b})\bigr{)}$ such that $\hat{s}^{b}_{i}(a)=(a,e_{i})$ for all $a$in $U_{b}$ (under the diffeomorphism $\Phi_{b}$). Since $\{U_{b}\}_{b\in B}$ is an open cover, by compactness, there exist points $b_{1},\ldots,b_{r}\in B$ such that $\{\,U_{b_{i}}\}_{i=1}^{r}$ remains an open cover of $B$. Let $\{\rho_{b_{i}}\}_{i=1}^{r}$ be a partition of unity subordinate to the open cover $\{\,U_{b_{i}}\}_{i=1}^{r}$, where each $\rho_{b_{i}}$ has compact support and admits a smooth square root. In this way, consider the global sections of $E^{V}$
@@
It is well–known that $\Gamma(E^{V})$ is isomorphic to the space of $G$–equivariant smooth functions $C^{\infty}_{\mathbb{C}}(P,V)^{G}=\{f:P\longrightarrow V\mid f\mbox{ is smooth such that }f(xA)=\alpha(A^{-1})f(x)\}$ as $\mathbb{C}^{\infty}_{\mathbb{C}}(B)$–bimodules. Then the maps
@@
where $\langle$-|$\rangle$ - | - ⟩ is the inner product that makes $\delta^{V}$ unitary (antilinear in the second coordinate). Now, let us consider the linear maps
@@
given by $T^{i}_{j}(e_{k})=f^{i}_{jk}$. A direct calculation shows that for all $x$in $P$ and for all $A$in $G$
@@ 4.1. Quantum Principal Bundles and Dunkl Operators
$\kappa:R\longrightarrow\mathbb{R}$ is a multiplicative function (it is $W(H)$–invariant), $r$in $R$ with $R$ the corresponding root system, and $r^{\#}$ is the element of the dual space associated with $r$. Covariant derivatives of this kind of qpc’s are given by
@@
for every $f$in $P$, which is a Dunkl operator in vector form [6].
Let $\delta^{V}$in $\mathcal{T}$. Then
@@
where $\mu^{D^{\omega}T}_{k}=\displaystyle\sum^{n_{V}}_{i=1}D^{\omega}T(e_{i})f^{V\,\ast}_{ki}$ and $T(e_{k})$in $P$ [1]. Notice that Dunkl connections are multiplicative, but not regular [6]. Furthermore, it is worth mentioning that, in differential geometry, there are no principal connections on the principal bundle $\rho:P_{\mathrm{class}}\longrightarrow P_{\mathrm{class}}/W(H)$.
@@
and evaluating on any $x$in $P_{\mathrm{class}}$ we have
@@
and evaluating on any $x$in $P_{\mathrm{class}}$ we get
@@
This implies for all $x$in $P_{\mathrm{class}}$ that $\langle\nabla^{\omega}_{V}(T_{1}),T_{2}\rangle_{\mathrm{L}}+\langle T_{1},\nabla^{\omega}_{V}(T_{2})\rangle_{\mathrm{L}}$ is equal to
@@
where $q=\displaystyle\sum^{n_{V}}_{k=1}T_{1}(e_{k})(\sigma_{r}\cdot T_{2}(e_{k}))^{\ast}$in $P$. However, since $T_{1}$, $T_{2}$in $\textsc{Mor}(\delta^{V},\Delta_{P})$, we have $\sigma_{r}\cdot T_{i}(e_{j})=\displaystyle\sum^{n_{V}}_{k=1}\lambda^{\sigma_{r}}_{kj}T_{i}(e_{k})$. Thus $q=\displaystyle\sum^{n_{V}}_{k=1}(\sigma_{r}\cdot T_{1}(e_{k}))T_{2}(e_{k})^{\ast}$ and hence $q(x)=q(x\sigma_{r})$ for all $x$in $P_{\mathrm{class}}$. In summary, we have proven explicitly that
@@
^$\langle\widehat{\nabla}^{\omega}_{V}(T_{1}),T_{2}\rangle_{\mathrm{R}}+\langle T_{1},\widehat{\nabla}^{\omega}_{V}(T_{2})\rangle_{\mathrm{R}}=d\langle T_{1},T_{2}\rangle_{\mathrm{R}}.$
@@
and by considering $\omega^{c}$ in equation (37), we have for all $\theta$in $\mathfrak{qg}^{\#}$
@@
where $\lambda$ is a Dunkl displacement, $\mu$in $\mathrm{Hor}^{1}P$, $x$in $P$, and $\theta$in $\mathfrak{qg}^{\#}$. This map is actually a qgt $\mathfrak{f}$ and a direct calculation shows that $\mathfrak{F}^{\circledast}\omega^{c}$ is a standard Dunkl connection. By using the form of the last map $\mathfrak{F}$, every single classical gauge transformation can be extended into a quantum one.

### v5

@@ 3. The Group Isomorphism
$\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})\otimes\mathbbm{1})\widetilde{\beta}((\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathfrak{F}^{-1})\mathrm{qtrs}(\vartheta^{(2)}))$
$\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})\otimes\hskip-19.91684pt\underbrace{\epsilon(\vartheta^{(2)})}_{\not=0\mathrm{\,only\,for\,}\Gamma^{\wedge\,0}=H}\hskip-19.91684pt\mathbbm{1})\,\widetilde{\beta}((\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathfrak{F}^{-1})\mathrm{qtrs}(\vartheta^{(3)}))$
$(-1)^{\partial\vartheta^{(2)}\partial[\vartheta^{(4)}]_{1}}\,\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\;\mathfrak{F}([\vartheta^{(1)}]_{2})\otimes\underbrace{\epsilon(\vartheta^{(2)})\mathbbm{1}}_{0-\mathrm{degree}})\,\widetilde{\beta}((\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathfrak{F}^{-1})\mathrm{qtrs}(\vartheta^{(3)})).$
$(-1)^{(\partial\vartheta^{(3)}+\partial\vartheta^{(2)})\partial[\vartheta^{(4)}]_{1}}\,\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})\otimes\underbrace{\vartheta^{(2)}S(\vartheta^{(3)}}_{0-\mathrm{degree}}))\,\widetilde{\beta}((\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathfrak{F}^{-1})\mathrm{qtrs}(\vartheta^{(4)}))$
$(-1)^{(\partial\vartheta^{(3)}+\partial\vartheta^{(2)})\partial[\vartheta^{(4)}]_{1}}\,\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})\otimes\underbrace{\vartheta^{(2)}S(\vartheta^{(3)}}_{0-\mathrm{degree}}))\,([\vartheta^{(4)}]_{1}\otimes\mathbbm{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}^{-1}([\vartheta^{(4)}]_{2}))$
$(-1)^{(\partial\vartheta^{(3)}+\partial\vartheta^{(2)})\partial[\vartheta^{(4)}]_{1}}\,\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})[\vartheta^{(4)}]_{1}\otimes\vartheta^{(2)}S(\vartheta^{(3)}))\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}^{-1}([\vartheta^{(4)}]_{2}))$
@@
$\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})[\vartheta^{(2)}]_{1}\otimes\mathbbm{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}^{-1}([\vartheta^{(2)}]_{2})),$
which proves our claim. Since $\mathfrak{F}^{-1}$ is a left $\Omega^{\bullet}(B)$–module morphism, by our previous claim and Proposition 3.3 point 11
@@
for every $\vartheta$ $\in$ $\Gamma^{\wedge}$. In a similar way, we can prove that $\mathfrak{f}_{\mathfrak{F}^{-1}}\ast\mathfrak{f}_{\mathfrak{F}}=\mathbbm{1}\epsilon$ and hence $\mathfrak{f}_{\mathfrak{F}}$ is a convolution invertible map.
@@
We are going to prove that $\mathfrak{F}_{\mathfrak{f}}$ is a graded left $\Omega^{\bullet}(B)$–module isomorphism which satisfies equation (70). First of all, by equation (68) it is obvious that $\mathfrak{F}_{\mathfrak{f}}(\mathbbm{1})=\mathbbm{1}$. Secondly, taking $\mu$ $\in$ $\Omega^{\bullet}(B)$ and $w$ $\in$ $\Omega^{\bullet}(P)$ we have $\mathfrak{F}_{\mathfrak{f}}(\mu w)=m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes\mathfrak{f})(\mu\otimes\mathbbm{1})\Delta_{\Omega^{\bullet}(P)}(w)=\mu\,\mathfrak{F}_{f}(w).$ Thirdly, by equation (68)
@@
is such that $\widehat{m}_{\Omega^{\bullet}}(w_{1}\otimes\vartheta_{1}\otimes w_{2}\otimes\vartheta_{2})=(-1)^{kl}w_{1}w_{2}\otimes\vartheta_{1}\vartheta_{2}$ if $w_{2}$ $\in$ $\Omega^{k}(P)$ and $\vartheta_{1}$ $\in$ $\Gamma^{\wedge l}$. In this way, by the graded Hopf $\ast$–algebra structure of $\Gamma^{\wedge\,\infty}$ and the fact that $\epsilon\not=0$ only for $\Gamma^{\wedge\,0}=H$, we get for all $w$ $\in$ $\Omega^{\bullet}(P)$
@@
Finally, notice that for all $w$ $\in$ $\Omega^{\bullet}(P)$
@@
are mutually inverse. Notice that for all $w$ $\in$ $\Omega^{\bullet}(P)$
@@
$=$ $(-1)^{\partial w^{(1)}\partial[w^{(2)}]_{1}}\,(w^{(0)}[w^{(2)}]_{1}\otimes\hskip-17.07182pt\underbrace{\epsilon(w^{(1)})}_{\not=0\mathrm{\,only\,for\,\Gamma^{\wedge\,0}=H}}\hskip-17.07182pt\mathbbm{1})$
@@
On the other hand, for every $\vartheta$ $\in$ $\Gamma^{\wedge}$
@@
for all $w$ $\in$ $\Omega^{\bullet}(P)$. Therefore $\mathfrak{f}_{1}\ast\mathfrak{f}_{2}\xmapsto{\widehat{\Lambda}^{-1}}\mathfrak{F}_{\mathfrak{f}_{1}}\circ\mathfrak{F}_{\mathfrak{f}_{2}}$ and the proposition follows. ∎
@@ 4. The form of the maps $\{T^{\mathrm{L}}_{k}\}$
Let $G$ be a compact matrix Lie group, and let $\mathcal{G}$ be its associated quantum group [2]. If $\delta^{V}$ is a (unitary) irreducible $\mathcal{G}$–corepresentation, then it induces a GG–representation on $V$. Indeed, by Theorem 1.1 we know that
@@
is a unitary and irreducible GG–representation. The following proposition is another of our purposes.
@@
Let $G$ be a compact matrix Lie group and let $\pi:P\longrightarrow B$ be a classical principal GG–bundle, where $P$ is the total space, $B$ is the base space and $\pi$ is the bundle projection. Assume $B$ are compact. If $\mathcal{T}$ is a complete set of mutually non–equivalent irreducible $\mathcal{G}$–corepresentations with $\delta^{\mathbb{C}}_{\mathrm{triv}}$ $\in$ $\mathcal{T}$, then for every $\delta^{V}$ $\in$ $\mathcal{T}$, there exists $\{T^{\mathrm{L}}_{k}\}^{d_{V}}_{k=1}$ $\subseteq$ $\textsc{Mor}(\delta^{V},\Delta_{P})$, for some $d_{V}$ $\in$ $\mathbb{N}$, such that equation (55) holds. Here
@@
is the pull–back of the right GG–action on $P$. In this setting, $C^{\infty}_{\mathbb{C}}(P)$ denotes the space of $\mathbb{C}$–valued smooth functions of $P$ and the tensor product is taken to be the completed injective tensor product.
@@
Let $\delta^{V}$ $\in$ $\mathcal{T}$ and consider its associated representation $\alpha^{V}$ with $n=\mathrm{dim}_{\mathbb{C}}(V)$. Since $\alpha^{V}$ is unitary, the associated vector bundle $\pi_{\alpha^{V}}:E^{V}\longrightarrow B$ is a Hermitian bundle.
For each $b$ $\in$ $B$, let $(U_{b},\Phi_{b})$ be a local trivialization of the Hermitian bundle $\pi_{\alpha^{V}}:E^{V}\longrightarrow B$ around $b$ associated to a principal GG–bundle local trivialization $(U_{b},\Psi_{b})$ of $\pi:P\longrightarrow B$. Then there exists a set of local sections $\{\hat{s}^{b}_{1},\ldots,\hat{s}^{b}_{1}\}$ $\subseteq$ $\Gamma\bigl(\pi_{{\alpha^{V}}}^{-1}(U_{b})\bigr)$ such that $\hat{s}^{b}_{i}(a)=(a,e_{i})$ for all $a$ $\in$ $U_{b}$ (under the diffeomorphism $\Phi_{b}$). Since $\{U_{b}\}_{b\in B}$ is an open cover, by compactness, there exist points $b_{1},\ldots,b_{r}\in B$ such that $\{\,U_{b_{i}}\}_{i=1}^{r}$ remains an open cover of $B$. Let $\{\rho_{b_{i}}\}_{i=1}^{r}$ be a partition of unity subordinate to the open cover $\{\,U_{b_{i}}\}_{i=1}^{r}$, where each $\rho_{b_{i}}$ has compact support and admits a smooth square root. In this way, consider the global sections of $E^{V}$
@@
It is well–known that $\Gamma(E^{V})$ is isomorphic to the space of GG–equivariant smooth functions $C^{\infty}_{\mathbb{C}}(P,V)^{G}=\{f:P\longrightarrow V\mid f\mbox{ is smooth such that }f(xA)=\alpha(A^{-1})f(x)\}$ as $\mathbb{C}^{\infty}_{\mathbb{C}}(B)$–bimodules. Then the maps
@@
where $\langle-|$-$\rangle$ is the inner product that makes $\delta^{V}$ unitary (antilinear in the second coordinate). Now, let us consider the linear maps
@@
given by $T^{i}_{j}(e_{k})=f^{i}_{jk}$. A direct calculation shows that for all $x$ $\in$ $P$ and for all AA $\in$ $G$
@@ 4.1. Quantum Principal Bundles and Dunkl Operators
$\kappa:R\longrightarrow\mathbb{R}$ is a multiplicative function (it is W(H)W(H)–invariant), $r$ $\in$ $R$ with $R$ the corresponding root system, and $r^{\#}$ is the element of the dual space associated with $r$. Covariant derivatives of this kind of qpc’s are given by
@@
for every $f$ $\in$ $P$, which is a Dunkl operator in vector form [6].
Let $\delta^{V}$ $\in$ $\mathcal{T}$. Then
@@
where $\mu^{D^{\omega}T}_{k}=\displaystyle\sum^{n_{V}}_{i=1}D^{\omega}T(e_{i})f^{V\,\ast}_{ki}$ and $T(e_{k})$ $\in$ $P$ [1]. Notice that Dunkl connections are multiplicative, but not regular [6]. Furthermore, it is worth mentioning that, in differential geometry, there are no principal connections on the principal bundle $\rho:P_{\mathrm{class}}\longrightarrow P_{\mathrm{class}}/W(H)$.
@@
and evaluating on any $x$ $\in$ $P_{\mathrm{class}}$ we have
@@
and evaluating on any $x$ $\in$ $P_{\mathrm{class}}$ we get
@@
This implies for all $x$ $\in$ $P_{\mathrm{class}}$ that $\langle\nabla^{\omega}_{V}(T_{1}),T_{2}\rangle_{\mathrm{L}}+\langle T_{1},\nabla^{\omega}_{V}(T_{2})\rangle_{\mathrm{L}}$ is equal to
@@
where $q=\displaystyle\sum^{n_{V}}_{k=1}T_{1}(e_{k})(\sigma_{r}\cdot T_{2}(e_{k}))^{\ast}$ $\in$ $P$. However, since $T_{1}$, $T_{2}$ $\in$ $\textsc{Mor}(\delta^{V},\Delta_{P})$, we have $\sigma_{r}\cdot T_{i}(e_{j})=\displaystyle\sum^{n_{V}}_{k=1}\lambda^{\sigma_{r}}_{kj}T_{i}(e_{k})$. Thus $q=\displaystyle\sum^{n_{V}}_{k=1}(\sigma_{r}\cdot T_{1}(e_{k}))T_{2}(e_{k})^{\ast}$ and hence $q(x)=q(x\sigma_{r})$ for all $x$ $\in$ $P_{\mathrm{class}}$. In summary, we have proven explicitly that
@@
$\langle\widehat{\nabla}^{\omega}_{V}(T_{1}),T_{2}\rangle_{\mathrm{R}}+\langle T_{1},\widehat{\nabla}^{\omega}_{V}(T_{2})\rangle_{\mathrm{R}}=d\langle T_{1},T_{2}\rangle_{\mathrm{R}}.$
@@
and by considering $\omega^{c}$ in equation (62), we have for all $\theta$ $\in$ $\mathfrak{qg}^{\#}$
@@
where $\lambda$ is a Dunkl displacement, $\mu$ $\in$ $\mathrm{Hor}^{1}P$, $x$ $\in$ $P$, and $\theta$ $\in$ $\mathfrak{qg}^{\#}$. This map is actually a qgt $\mathfrak{f}$ and a direct calculation shows that $\mathfrak{F}^{\circledast}\omega^{c}$ is a standard Dunkl connection. By using the form of the last map $\mathfrak{F}$, every single classical gauge transformation can be extended into a quantum one.